OpenStudy (anonymous):

3.Your piggy bank has a total of 46 coins in it;some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) 3.Your piggy bank has a total of 46 coins in it;some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) @Mathematics

OpenStudy (anonymous):

Let D be the number of dimes, Q quarters. Total number of coins: D+Q=46......................(1) Total value of coins: 10D+25Q=700 2D+5Q=140.......................(2) From eq. 1, D=46-Q Substitute in eq. 2: 92-2Q+5Q=140 3Q=48 Q=16 ...then : D=46-Q ...D=30 Give me a Medal plz (be fan of me )

OpenStudy (anonymous):

you wanna do a algebraching essay for me ?

OpenStudy (anonymous):

What did u mean by (algebraching essay)?? didn't get it ..srry

OpenStudy (anonymous):

1.Find a partner. 2.In Paint, PowerPoint, a program of your choosing, or using a scan of a handmade drawing, graph a system of inequalities on a coordinate plane. Please limit your coordinate plane to a 10 × 10 grid. The x-axis and y-axis should go from –5 to 5. Write the system of inequalities that accompanies your graph. 3.Pick an ordered pair within the shaded region to bury your “treasure.” Keep this ordered pair to yourself and do not share it with your partner. 4.Submit the system of inequalities in slope-intercept form to your partner. You will not share your graph or the hidden treasure point, only the system of inequalities. Since your family member or friend is not in this course you will need to supply them with step by step directions on how to graph this system of inequalities. If your partner does not successfully guess the ordered pair you may want to take a look at their graph. At this point, you may give your partner helpful hints on how they may fix their graph or narrow the search by giving them a hint. Your hint may be an equation of another line that the treasure point lies on. Record all of the guesses it takes your partner to find where your treasure point lies. Whether you are doing this activity with a family member or friend, or collaboratively with another student, you will submit the following to your instructor as assignment “06.07 Algebracaching": 1.Your system of inequalities, the graph, and hidden treasure point. 2.The step by step directions you provided your partner to graph the system of inequalities. 3.The guesses your partner tried to find your treasure point. 4.Your reflection — write a paragraph, of 3 or more sentences, as you reflect on this activity and your final product. You may use specific questions to guide your thoughts, if needed.

OpenStudy (anonymous):

your my partner. i dont know exactly what to do....

OpenStudy (anonymous):

That's really too long , It requires a long solution !I can't give it to u here .I can send it to u via email!

OpenStudy (anonymous):

will you send it to me within the next hr and a half ? if it might take you that long ?

OpenStudy (anonymous):

copy quickly

OpenStudy (anonymous):

whats ur email

OpenStudy (anonymous):

dont see u... add me

OpenStudy (anonymous):

i am there :D

OpenStudy (anonymous):

message me

OpenStudy (anonymous):

can u see me

OpenStudy (anonymous):

no

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i emailed u

OpenStudy (anonymous):

im not scared i just need my assignment done....

OpenStudy (anonymous):

i dont wanna flirt.. i wanna graduate highschool

OpenStudy (anonymous):

I wasn't flirting :S . If I say 'babe' ,I mean u r my mate ,,,,my close mate ..srry for offence :(

OpenStudy (anonymous):

lol ok its fine... ur from doha ?

OpenStudy (anonymous):

yup =) ..how old r u? .I am older than u by the way :D .Not gonna flirt with my sweet young sister!

OpenStudy (anonymous):

lol seen that and in almost a month i'll be 17... yeah i am way younger.... & sister ? ok i'll play that game,, brother, will you do my assignment ?

OpenStudy (anonymous):

lol..where r u from?

OpenStudy (anonymous):

fl.

OpenStudy (anonymous):

finland?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Oh..The place where Nokia was born :D

OpenStudy (anonymous):

i guess... lol will you do this assignment plz... its my last one

OpenStudy (anonymous):

I will but not now .It gonna take a lot of time :S

OpenStudy (anonymous):

u should try solving it girlll..don't be lazyyy

OpenStudy (anonymous):

im not its just fricken hard

OpenStudy (anonymous):

i sent it in the msn it finally works

OpenStudy (anonymous):

Not to be rude, but what in the world are we supposed to do to help you out here? This is all about drawing, partnering with a friend, and playing what appears to be algebra Battleship, all of which are darned near impossible to do in a text box on Y!A.

OpenStudy (anonymous):

wel, all i need you to do is do it in a paint file or something and send it in an email with everything needed and i will draw what you did and send it in my own words

OpenStudy (anonymous):

Sorry..I can't help u in this :(